OpenStudy (anonymous):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round. Which equation can be used to represent this situation? A. p – (–8) = –12.5 B. p + 12.5 = –8 C. p – 12.5 = –8 D. p + 8 = 12.5

OpenStudy (ilovecake):

p – 12.5 = –8

OpenStudy (ilovecake):

ugh

OpenStudy (ilovecake):

It is C

OpenStudy (anonymous):

ok how do you know?

OpenStudy (ilovecake):

Okay. Because I did this before and I got 100%

OpenStudy (anonymous):

ok

OpenStudy (ilovecake):

I swear it is C

OpenStudy (anonymous):

the last one is A zoologist is recording the loss of wolves in her state. She notes the number of wolves, w, in the state on January 1. One year later, there were 84 wolves in the state, which is 23 fewer wolves than were in the state a year earlier. Which equation represents this situation? A. w + 23 = 84 B. w – 23 = 84 C. 23 – w = 84 D. w – 84 = –23

OpenStudy (ilovecake):

w+23=84. I remember that one

OpenStudy (anonymous):

so a???

OpenStudy (ilovecake):

yeah. :)

OpenStudy (anonymous):

ok hold on i will tell you what i got

OpenStudy (ilovecake):

I hope you pass. :)

OpenStudy (anonymous):

ok so i got 40% but all of your answers where correct

OpenStudy (anonymous):

GJ good job

OpenStudy (ilovecake):

How many questions did you have?

OpenStudy (anonymous):

5

OpenStudy (ilovecake):

then how did you get a 40% when my answers were right?

OpenStudy (anonymous):

i do not know???? hold on

OpenStudy (ilovecake):

I am sooo sorry :(

OpenStudy (ilovecake):

if you need help again msg me. :)

OpenStudy (anonymous):

ok

OpenStudy (ilovecake):

bye.